OpenStudy (anonymous):

What is the equation in standard form of a perpendicular line that passes through (5,1)?

OpenStudy (anonymous):

Hello ,Check this link . It will help you a lot !! ( http://openstudy.com/study#/updates/4ff5ff4de4b01c7be8c90c78)

OpenStudy (phi):

perpendicular to what ? They must tell you another line to be perpendicular to. Find the slope of that line (if they didn't tell you). If we call its slope m, the slope of the perpendicular line will be - 1/m (if m is a fraction, "flip it") example: m=1/2 then - 1/m would be - 2/1 or just -2 Finally use y - y0 = -1/m ( x - x0) with (x0,y0)= (5,1)

OpenStudy (anonymous):

I am not given the slope so how do I find it with just those two points?

OpenStudy (anonymous):

I need an equation for a perpendicular line that passes through the points (5,1).

OpenStudy (anonymous):

I am really confused because I am not given any more information, so can anybody help me?

OpenStudy (phi):

perpendicular means two lines that cross and form a right angle (90 degree angle) we can make up lots of lines that are perpendicular to each other where one (or both ) go through point (5,1). Can you post a screen shot of the exact question ?

OpenStudy (anonymous):

Aden are you Asian?

OpenStudy (anonymous):

Another question follows it: What is the x-intercept of the perpendicular line?

OpenStudy (anonymous):

I'm supposed to find the answer with just the two points (5,1).

OpenStudy (anonymous):

don't know how to do that...

OpenStudy (tkhunny):

There is some possible language confusion. In at least some Chinese, Korean, and Japanese dialects, it appears there may be some ambiguity between the English "vertical" and the English "perpendicular". Thus, "perpendicular", when translated may mean "perpendicular to the horizon", or "vertical". If the point is (5,1), the horizontal line through that point is y = 1 If the point is (5,1), the vertical ("perpendicular") line through that point is x = 5

OpenStudy (anonymous):

how will this information help me?

OpenStudy (anonymous):

I am not Asian btw.

OpenStudy (anonymous):

I got y=1/mx+1+5m

OpenStudy (anonymous):

When I changed it to standard form, I got my=x+m+5

OpenStudy (anonymous):

I mean my=-x+m+5

OpenStudy (anonymous):

How do I get the x intercept of the perpendicular line now?

OpenStudy (tkhunny):

Nevertheless. If the problem statement truly says, simple "perpendicular line through a point", the question makes no sense at all unless we can make it make sense. I find the same ambiguity in some eastern and western languages. In the standard American English, "find the perpendicular line through (5,1)" makes no sense and the problem cannot be solved. If we assume that "perpendicular" has been translated from some language where it is culturally assumed that "perpendicular" means "perpendicular to the horizon" unless otherwise defined, then the question can make sense. In standard English, it would be "find the vertical line through (5,1)" and the answer is x = 5. Your answer makes no sense at all. How did you get that?

OpenStudy (tkhunny):

If you are going to speak standard English mathematics, you have to quit saying "perpendicular line". It does not mean anything without reference to a given, additional line. It has to be perpendicular to something. It can't be simply perpendicular.

OpenStudy (anonymous):

Oh okay. I thought I know you

OpenStudy (anonymous):

I really don't know...

OpenStudy (anonymous):

I have been working on this question for days now...

OpenStudy (anonymous):

I don't know how they expect me to find the answer without giving what line it is perpendicular to, y'know.

OpenStudy (tkhunny):

Well, stop it. I have provided as thorough and complete an answer as can be provided with the given information. If you still are not getting it, you need to have a good talk with your teacher/advisor/author. It is possible the question has been written incorrectly. That should not be a problem. Point it out and move on.

OpenStudy (anonymous):

That's what I should do, you're right. It could be written incorrectly. I believe it is. Thank you for your help so far!